A truck travelling at 54 km/h is slow down to 36 km/h in 10 sec. Find the retardation


1
Expert's answer
2021-06-09T17:54:00-0400
"a=\frac{v_i-v_f}{t}\\a=\frac{54-36}{3.6\cdot10}=0.5\frac{m}{s}"